1) 3x=7y+2;(2,3)
2) 11=3y+2x

Answer by 332915(1)

1) -7y=-3x+2
y=3/7x+2
m=3/7
b=2
y-y1=m(x-x1)
y-3=3/7*(x-2)
y-3=3/7x-6/7
y=3/7x+2 1/7



2) -3y=2x-11
y=-2/3x+11/3
m=-2/3
b=11/3
y-y1=m(x-x1)
y-3=2/3*(x-2)
y-3=2/3x-4/3
y=2/3x+2 1/3